John is 35 years old, which is 5 years less than b times Bob's age. How old is Bob?
Found 2 solutions by FrankM, fractalier:
Answer by FrankM(1040)

35 = bB-5
bB=40
B=40/b

Answer by fractalier(6550)

Let's call Bob's age, x. Then we can write
35 = bx - 5
Now we solve for x...we add 5 first and get
40 = bx
Now divide by b and get
x = 40/b
We cannot know Bob's age exactly until we know b...